OpenStudy (rainbow_dashie):

Solve for x. @satellite73

OpenStudy (rainbow_dashie):

\[_{3}(-4x-6)=_{3}(-6x+2)\]

OpenStudy (anonymous):

wht is that little 3 doing there ?

OpenStudy (anonymous):

\[3^{-4x-6}=3^{-6x+2}\]?

OpenStudy (rainbow_dashie):

its just a 3

OpenStudy (anonymous):

\[3(-4x-6)=3(-6x+2)\]?

OpenStudy (rainbow_dashie):

yes

OpenStudy (anonymous):

then ignore the 3 since it is on both sides and solve \[-4x-6=-6x+2\] in 3 steps

OpenStudy (anonymous):

1) add \(6x\) to both sides 2) add \(6\) to both sides 3) divide both sides by \(2\)

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

i see you changed your picture!

OpenStudy (anonymous):